An airplane leaves an airport and flies due west 120 miles and then 150 miles in the direction S 39.17°W. How far is the plane from the airport (round to the nearest mile)?
Answer by kieran(1)

x = 120 * cos(0) + 150 * cos(-39.17)
y = 120 * sin(0) + 150 * sin(-39.17)
x = 120 + 150 * cos(39.17)
y = -150 * sin(39.17)
sqrt(x^2 + y^2) =>
sqrt(14400 + 36000 * cos(39.17) + 22500 * cos(39.17)^2 + 22500 * sin(39.17)^2) =>
sqrt(14400 + 36000 * cos(39.17) + 22500) =>
sqrt(36900 + 36000 * cos(39.17)) =>
sqrt(900 * (41 + 40 * cos(39.17))) =>
30 * sqrt(41 + 40 * cos(39.17)) =>
244